Calculate the economic order quantity for the scenario above, if annual demand is 3080 jobs The opportunity cost of money is = 15% per year. The cost physically hold an item in inventory is 8% per year. The cost to the company of ALL the material for one job is $750 Administrative Order cost of $100.

What is the EOQ?
Answer #1

Given information :

Annual Demand (D) =3080 Jobs

Opportunity cost of money=15% per year

Inventory holding cost per year =8%

Material cost for one job =$750

Administrative order cost (S) =$100

Material cost per year = $ Annual demand * Material cost for one job =$ 3080*750 =$23,10,000

Inventory holding cost per year =$23,10,000 * 8/100 =$ 1,84,800

Inventory holding cost/unit/year (H) =1,84,800/3080 =$60

We have the formula for EOQ

EOQ =Sq Root (2SD/H)=Sq.Root ( 2*100*3080/60)=Sq.Root ( 10266.66)=101.32 =101 (Round Fig)

EOQ = 101 Jobs

Please note that EOQ is independent of oppurtunty cost .That means whtever be the oppurtunity cost economic order quantity will not change.